# PortLLM: Personalizing Evolving Large Language Models with Training-Free and Portable Model Patches

### v2

@@ 4.1 Superiority of PortLLM Framework
In this section, we compare the performance of our model patches against several baseline models, including the pretrained LLM $\theta$, the updated model using continued pretraining $\theta^{\prime}$, the fine-tuned model $\theta_{i}$, and the updated fine-tuned model $\theta^{\prime}_{i}$. For consistency, all models are variations of Mistral-7B, while continued pretraining dataset is OpenOrca. Importantly, the performance is evaluated under zero-shot setting across all the datasets. The results are summarized in Table 2.
@@
❷ Comparing the performance of our model patches applied to the updated model $\theta^{\prime}+\Delta\theta_{i}$ with that of the fine-tuned model $\theta^{\prime}$, we observe that our method successfully transfers most of the downstream task-specific knowledge, yielding comparable results. As shown in Table 2, for tasks {BoolQ, MRPC, WinoGrande, WNLI}, the performance is nearly identical, with a maximum difference of only $0.77\%$ in favor of the fine-tuned model. However, in tasks like {SST-2, RTE, GSM8K}, our approach outperforms the fine-tuned model by {$0.11\%,1.44\%,7.28\%$}, respectively. These results suggest that, when paired with a pretraining dataset that enhances performance for a specific task, our model patches can further leverage this advantage to improve downstream task performance in certain cases.
❸ Moreoever, our method performs on par with the fine-tuned updated model ($\theta^{\prime}+\Delta\theta_{i}$ compared to $\theta_{i}^{\prime}$), as evident from the comparison of the last two rows in Table 2. The difference between the two approaches are minimal when it comes to performance, with a maximum variation of just $1.04\%$ observed in the case of WinoGrande. Notably, while one method requires fine-tuning, our approach remains completely training-free. Additionally, for certain downstream tasks such as WNLI and GSM8K, our method outperforms the fine-tuned updated model by $1.02\%$ and $6.37\%$, respectively. This demonstrates that our approach not only provides comparable results but can, in some instances, surpass the performance of a fine-tuned evolved model.
4.2 Consistent Results across Different Pretraining Datasets
@@
4.3 Consistent Results across Different Model Architectures
This section provides a comprehensive analysis of our model patches across various architectures to evaluate their performance. We examine four different model architectures: {Mistral-7B, Llama2-7B, Llama3.1-8B, Gemma2-9B}, assessing their effectiveness on four distinct downstream tasks {BoolQ, MPRC, WNLI, WinoGrande}. Performance is evaluated under four settings : (1) pretrained model, (2) continued pretrained model or Updated Model, (3) fine-tuned model, and (4) our model patches ported to the updated model. The OpenOrca dataset is used for our continued pretraining in this analysis. The results are summarized in Table 4
❶ The results across different model architectures indicate that our training-free model patches significantly enhance zero-shot performance for all downstream tasks. In each case, our approach matches personalized performance (fine-tuned model), and in certain instances, it even surpasses it. For example, with Gemma2-9B, when the personalized performance exceeds our method, the difference in accuracy is at most $0.73\%$, which can considered negligible. Conversely, in scenarios where our method outperforms personalized performance, we observe improvements of up to $5.53\%$. A similar trend is noted across the other model architectures, as detailed in Table 4.
Table 4: Performance analysis of model patches across various architectures {Mistral-7B, Llama2-7B, Llama3.1-8B, Gemma2-9B} on four downstream tasks {BoolQ, MRPC, WNLI, WinoGrande} under four different model settings. For each downstream task bold highlights the best performance for each model architecture.
@@ 4.3 Consistent Results across Different Model Architectures
Fine-tuned Model $\theta_{i}$ 91.38 91.42///93.83 $83.20$ 83.98
@@ 4.4 PortLLM Also Works with Full Weight Continued Pretraining
We find that our model patches can be effectively applied to a continued pretrained model utilizing full weight updates, rather than relying solely on LoRA. Across all evaluated datasets – {BoolQ, SST-2, MRPC, RTE, WinoGrande, WNLI, GSM8K} – we observe significant performance improvements of $\{3.27\%,1.72\%,10.04\%,15.52\%,12.31\%,16.90\%,20.69\%\}$, respectively, compared to the zero-shot performance of the Updated Model.
Table 6: Efficiency Comparison Between PortLLM and LoRA on SST-2 with Mistral-7B as the model architecture. The table compares trainable parameters, GPU Memory Usage, and GPU Hours for PortLLM, and LoRA fine-tuning.
@@ 4.5 Computing Efficiency Comparison of PortLLM
This section evaluates the performance of our method from an efficiency perspective. We employ the following metrics for comparison: (1) Number of Trainable Parameters, (2) GPU Memory Utilization, and (3) GPU Hours. We analyze the merging of our model patches in relation to model fine-tuning using LoRA to achieve comparable performance. For LoRA fine-tuning calculations, we have the following settings: Downstream task SST-2 for Mistral-7B, with local Batch Size of $4$ and $5$ Epochs. The results are summarized in Table 6.
Compared to downstream fine-tuning using LoRA, our methods offers plug-and-play solution with no trainable parameters, resulting in a reduction of nearly $20$ million parameters that need to be trained. This training-free paradigm not only conserves resources but also saves up to 12.2×12.2$\times$12.2 × GPU Memory, reducing the requirement from $350$ GB for LoRA to just $28.7$ GB during the merge operation of model patches. Additionally, the merge operation can be executed in mere seconds, in start contrast to the hours required for fine-tuning. This opens many other doors for applications of our model patches. PortLLM demonstrates the potential for on-device, training-free models for various down-stream tasks without the need for fine-tuning. Furthermore, it reduces the need for expensive cloud infrastructure, especially in large-scale fine-tuning.
@@ 5 Conclusion
In this paper, we propose PortLLM, a framework aimed at addressing the challenges faced by downstream users of pretrained LLMs when adapting to frequent model evolutions over time. By leveraging lightweight model patches, PortLLM offers a training-free, cost-effective solution to seamlessly transfer domain-specific knowledge between different iterations of LLMs. This enables users to maintain, and sometimes even enhance, their models’ performance on specialized tasks without the need for repeated fine-tuning or extensive computational resources. Through extensive empirical evaluations across a set of tasks and models, we demonstrate that our method not only preserves performance but can also leverage the continual updates in pretraining LLMs, offering substantial gains in-task specific performance. Moreover, we provide theoretical insights into the portability of these model patches, highlighting the underlying factors that make them effective across evolving model versions. Looking forward, PortLLM paves the way for more robust and adaptable solutions in the evolving landscape of LLM personalization, offering another avenue for training-free adaptation. Furthermore, future endeavours will aim at developing such methods that work across different model architecture, including using techniques from model merging.

### v3

@@ 4.1 Superiority of PortLLM Framework
In this section, we compare the performance of our model patches against several baseline models, including the pretrained LLM $\theta$, the updated model using continued pretraining $\theta^{\prime}$, the fine-tuned model $\theta_{i}$, and the updated fine-tuned model $\theta^{\prime}_{i}$. For consistency, all models are variations of Mistral-7B, while the continued pretraining dataset is OpenOrca. Importantly, the performance is evaluated under the zero-shot setting across all the datasets. The results are summarized in Table 2.
@@
❷ Comparing the performance of our model patches applied to the updated model $\theta^{\prime}+\Delta\theta_{i}$ with that of the fine-tuned model $\theta_{i}$, we observe that our method successfully transfers most of the downstream task-specific knowledge, yielding comparable results. As shown in Table 2, for tasks {BoolQ, MRPC, WinoGrande, WNLI}, the performance is nearly identical, with a maximum difference of only $0.77\%$ in favor of the fine-tuned model. However, in tasks like {SST-2, RTE, GSM8K}, our approach outperforms the fine-tuned model by {$0.11\%,1.44\%,7.28\%$}, respectively. These results suggest that, when paired with a pretraining dataset that enhances performance for a specific task, our model patches can further leverage this advantage to improve downstream task performance in certain cases.
❸ Moreover, our method performs on par with the fine-tuned updated model ($\theta^{\prime}+\Delta\theta_{i}$ compared to $\theta_{i}^{\prime}$), as evident from the comparison of the last two rows in Table 2. The difference between the two approaches is minimal when it comes to performance, with a maximum variation of just $1.04\%$ observed in the case of WinoGrande. Notably, while one method requires fine-tuning, our approach remains completely training-free. Additionally, for certain downstream tasks such as WNLI and GSM8K, our method outperforms the fine-tuned updated model by $1.02\%$ and $6.37\%$, respectively. This demonstrates that our approach not only provides comparable results but can, in some instances, surpass the performance of a fine-tuned evolved model.
@@
4.2 Consistent Results across Different Pretraining Datasets
@@
Table 4: Performance analysis of model patches across various architectures {Mistral-7B, Llama2-7B, Llama3.1-8B, Gemma2-9B} on four downstream tasks {BoolQ, MRPC, WNLI, WinoGrande} under four different model settings. For each downstream task Cyan highlights the best performance for each model architecture.
@@
Fine-tuned Model $\theta_{i}$ $91.38$ $91.42$ / $93.83$ $83.20$ $83.98$
@@
4.3 Consistent Results across Different Model Architectures
This section provides a comprehensive analysis of our model patches across various architectures to evaluate their performance. We examine four different model architectures: {Mistral-7B, Llama2-7B, Llama3.1-8B, Gemma2-9B}, assessing their effectiveness on four distinct downstream tasks {BoolQ, MPRC, WNLI, WinoGrande}. Performance is evaluated under four settings: (1) pretrained model, (2) continued pretrained model or updated model, (3) fine-tuned model, and (4) our model patches ported to the updated model. The OpenOrca dataset is used for our continued pretraining in this analysis. The results are summarized in Table 4.
❶ The results across different model architectures indicate that our training-free model patches significantly enhance zero-shot performance for all downstream tasks. In each case, our approach matches personalized performance (fine-tuned model), and in certain instances, it even surpasses it. For example, with Gemma2-9B, when the personalized performance exceeds our method, the difference in accuracy is at most $0.73\%$, which can be considered negligible. Conversely, in scenarios where our method outperforms personalized performance, we observe improvements of up to $5.53\%$. A similar trend is noted across the other model architectures, as detailed in Table 4.
@@ 4.4 PortLLM Also Works with Full Weight Continued Pretraining
We find that our model patches can be effectively applied to a continued pretrained model utilizing full weight updates rather than relying solely on LoRA. Across all evaluated datasets – {BoolQ, SST-2, MRPC, RTE, WinoGrande, WNLI, GSM8K} – we observe significant performance improvements of $\{3.27\%,1.72\%,10.04\%,15.52\%,12.31\%,16.90\%,20.69\%\}$, respectively, compared to the zero-shot performance of the updated model.
Table 6: Efficiency comparison between PortLLM and LoRA on SST-2 with Mistral-7B as the model architecture. The table compares trainable parameters, GPU Memory Usage, and GPU Hours for PortLLM and LoRA fine-tuning.
@@ 4.5 Computing Efficiency Comparison of PortLLM
This subsection evaluates the performance of our method from an efficiency perspective. We employ the following metrics for comparison: (1) Number of trainable parameters, (2) GPU memory utilization, and (3) GPU hours. We analyze the merging of our model patches in relation to model fine-tuning using LoRA to achieve comparable performance. For LoRA fine-tuning calculations, we have the following settings: Downstream task SST-2 for Mistral-7B, with local batch size of $4$ and $5$ epochs. The results are summarized in Table 6.
Compared to downstream fine-tuning using LoRA, our method offers a plug-and-play solution with no trainable parameters, resulting in a reduction of nearly $20$ million in parameters that need to be trained. This training-free paradigm not only conserves resources but also saves up to 12.2×12.2$\times$12.2 × GPU memory, reducing the requirement from $350$ GB for LoRA to just $28.7$ GB during the merge operation of model patches. Additionally, the merge operation can be executed in mere seconds, in contrast to the hours required for fine-tuning. This opens many other doors for applications of our model patches. PortLLM demonstrates the potential for on-device, training-free models for various downstream tasks without the need for fine-tuning. Furthermore, it reduces the need for expensive cloud infrastructure, especially in large-scale fine-tuning.
@@ 5 Conclusion
In this paper, we propose PortLLM, a framework aimed at addressing the challenges faced by downstream users of pretrained LLMs when adapting to frequent model evolutions over time. By leveraging lightweight model patches, PortLLM offers a training-free, cost-effective solution to seamlessly transfer domain-specific knowledge between different iterations of LLMs. This enables users to maintain, and sometimes even enhance, their models’ performance on specialized tasks without the need for repeated fine-tuning or extensive computational resources. Through extensive empirical evaluations across a set of tasks and models, we demonstrate that our method not only preserves performance but can also leverage the continual updates in pretrained LLMs, offering substantial gains in task-specific performance. Moreover, we provide theoretical insights into the portability of these model patches, highlighting the underlying factors that make them effective across evolving model versions. Looking forward, PortLLM paves the way for more robust and adaptable solutions in the evolving landscape of LLM personalization, offering another avenue for training-free adaptation. Furthermore, future endeavours will aim at developing such methods that work across different model architecture, including using techniques from model merging.
6 Reproducibility Statement
To ensure reproducibility, we provide detailed descriptions of datasets, model architectures, training settings, and evaluation metrics used in our experiments in Section 4. The same section also dives deep into the hyperparameters used for all the tasks mentioned in our paper, including LoRA fine-tuning and downstream evaluation. Furthermore, we have also provided all the training scripts alongside the hyperparameters as supplementary material so that results from our papers can be reproduced with minimal effort. Lastly, the datasets and model architectures utilized in this paper are open-source and publicly available for anyone’s use. Each dataset, as well as model, have been cited accordingly so that anyone can reproduce the experiments.
Acknowledgment
This research was, in part, funded by the CISCO Faculty Award, UNC SDS Seed Grant and NetMind.AI. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing official policies, either expressed or implied of the funding organizations.
Appendix
Appendix A Analysis of Model Performance Under Multiple Continual Updates
To validate our framework’s robustness under periodic updates, we conducted experiments simulating multiple rounds of continued pretraining. Using Mistral-7B as our base model, we performed four sequential updates using different pretraining datasets: OpenOrca $\to$ OpenPlatypus $\to$ Alpaca $\to$ GPT4-LLM-Cleaned. For the hyperparameters, we utilize the same settings described in Section 4. We evaluated our model patch on all seven downstream tasks after each update. The results are summarized in Table A. Our experiments yield the following key findings:
Table A: Model performance across sequential updates ($T=0$ to $T=4$) on seven downstream tasks. We report accuracy for all tasks except MRPC, where we show both accuracy and F1 score.
Time Model Version BoolQ SST-2 MRPC RTE WinoGrande WNLI GSM8K
Accuracy Accuracy Accuracy/F1 Accuracy Accuracy Accuracy Accuracy
$T=0$ Pretrained Model $\theta$ $83.58$ $66.86$ $65.20$ / $73.70$ $67.51$ $74.11$ $57.76$ 6.37
None Fine-tuned Model $\theta_{i}$ 91.01 95.99 89.46 / 92.62 87.73 85.95 83.11 34.04
$T=1$ Updated Model $\theta^{\prime}$ 87.46 82.91 74.75 / 83.73 75.09 75.06 57.72 15.16
OpenOrca Ours $\theta^{\prime}+\Delta\theta_{i}$ 90.24 96.11 88.73 / 92.10 89.17 85.01 83.10 41.32
$T=2$ Updated Model $\theta^{\prime}$ 86.33 86.81 76.23 / 83.53 71.48 74.51 52.11 12.36
OpenPlatypus Ours $\theta^{\prime}+\Delta\theta_{i}$ 89.88 96.22 88.24 / 91.55 88.09 84.37 83.10 42.15
$T=3$: Updated Model $\theta^{\prime}$ 87.03 85.78 74.02 / 83.33 73.65 74.27 56.34 16.38
Alpaca Ours $\theta^{\prime}+\Delta\theta_{i}$ 89.66 96.33 88.73 / 92.12 88.81 85.08 83.10 38.21
$T=4$: Updated Model $\theta^{\prime}$ 85.11 75.79 72.78 / 82.79 74.37 71.67 56.34 14.94
GPT4-LLM Ours $\theta^{\prime}+\Delta\theta_{i}$ 88.41 96.10 87.01 / 90.91 85.56 80.19 77.46 31.77
❶ Across all update stages ($T=0$ to $T=4$), applying our model patches results in substantial zero-shot improvements over the updated model $\theta^{\prime}$. These improvements are consistent and significant across different tasks, with SST-2 showing gains from $+9.41\%$ to $+20.31\%$, WNLI maintaining strong improvements between $+21.12\%$ and $+30.99\%$, and MRPC consistently improving by $+12$ to $14\%$ in accuracy. Notably, these improvements are achieved without any additional training, requiring only a simple merge operation of our model patches.
❷ The performance stability of our patched models is particularly noteworthy when compared to the fluctuating zero-shot performance of $\theta^{\prime}$. As shown in Table A, our method maintains remarkably consistent performance across multiple updates. For instance, BoolQ accuracy remains within a tight range of $88.41$ to $90.24\%$, SST-2 consistently maintains accuracy above $96\%$, and MRPC’s F1 score stays above $90$ across all update stages. These results demonstrate our method’s robustness to successive model updates and its ability to preserve task-specific knowledge.
❸ Furthermore, our method shows interesting behavior in leveraging complementary knowledge from different updates. Taking GSM8K as an example, we observe varying but significant improvements ranging from $+16.83\%$ to $+29.79\%$ across different update stages. This suggests that our model patches can effectively combine knowledge from both the original fine-tuning and the continued pretraining updates, sometimes leading to performance gains that exceed what might be expected from either source alone. Such behavior demonstrates the potential of our approach to not just preserve but potentially enhance task performance through knowledge integration across model versions.
Appendix B Analysis of LoRA Rank Selection for Downstream Tasks
To validate our choice of LoRA rank and understand its impact on model performance, we conducted experiments with varying ranks. This analysis helps establish the optimal balance between computational efficiency and model effectiveness. To perform these experiments, we utilize Mistral-7B with OpenOrca as the continued pretraining dataset and evaluate on BoolQ, MRPC, and WNLI downstream tasks with a varying rank for LoRA. The rest of the hyperparameters are the same as in Section 4. The results are shown in Table B.
❶ Both fine-tuning $\theta_{i}$ and our method $\theta^{\prime}+\Delta\theta_{i}$ demonstrate remarkable stability across different ranks, with optimal performance typically achieved at rank 8 to 16. Specifically, on BoolQ, we observe peak accuracy at rank 16, while MRPC and WNLI show optimal performance at rank 8. This consistency across ranks validates the robustness of our approach and suggests that model patches effectively capture task-specific knowledge regardless of rank selection.
❷ When examining the efficiency aspects, we find that increasing rank beyond 8 provides diminishing or even negative returns. As shown in Table B, at rank 32, we observe decreased performance across most tasks compared to rank 8: BoolQ drops by $1.01\%$, MRPC by $0.49\%$, and WNLI by $2.83\%$. Importantly, the performance gap between our method and direct fine-tuning remains minimal even at lower ranks, with differences of less than $1\%$ in most cases. This suggests that our training-free approach maintains its effectiveness even with more constrained rank settings.
❸ Our analysis strongly justifies our original choice of rank 8 as the default setting. This configuration achieves an optimal balance between computational efficiency (requiring fewer parameters than higher ranks), model performance (maintaining competitive results across all tasks), and adaptation capability (providing sufficient capacity for task-specific learning). Notably, while higher ranks like 16 or 32 require significantly more parameters, they offer minimal or no performance benefits, making rank 8 the sweet spot for our training-free framework.
Table B: Performance comparison across different LoRA ranks (2, 4, 8, 16, 32) on three downstream tasks using Mistral-7B.
Rank Model Version BoolQ MRPC WNLI
Accuracy Accuracy/F1 Accuracy
$r=2$ Fine-tuned Model $\theta_{i}$ 90.52 87.30 / 91.28 74.65
Ours $\theta^{\prime}+\Delta\theta_{i}$ 89.85 87.01 / 91.03 77.46
$r=4$ Fine-tuned Model $\theta_{i}$ 90.81 89.42 / 91.98 78.23
Ours $\theta^{\prime}+\Delta\theta_{i}$ 90.18 87.75 / 91.53 80.28
$r=8$ Fine-tuned Model $\theta_{i}$ 91.01 89.46 / 92.62 83.11
Ours $\theta^{\prime}+\Delta\theta_{i}$ 90.24 88.73 / 92.10 83.10
$r=16$ Fine-tuned Model $\theta_{i}$ 91.07 89.22 / 92.49 81.69
Ours $\theta^{\prime}+\Delta\theta_{i}$ 90.98 88.97 / 92.31 82.98
$r=32$ Fine-tuned Model $\theta_{i}$ 90.00 88.97 / 92.15 80.28
Ours $\theta^{\prime}+\Delta\theta_{i}$ 90.28 88.73 / 91.93 81.69
Appendix C Proof of Lemma 1
Notations: $\mathcal{C}(\cdot)$ returns the column vector subspace of a matrix.
Recall in Section 3.3, we decomposed the fine-tuned updated model $\theta^{\prime}_{i}$ into two terms:
$\theta_{i}^{\prime}$ $=\theta^{\prime}+\Delta\theta_{i}^{\prime}$ (8a)
$=\underbrace{(\theta^{\prime}+\Delta\theta_{i})}_{\text{Naive Update }\hat{\theta}_{i}^{\prime}}+\underbrace{(\Delta\theta_{i}^{\prime}-\Delta\theta_{i})}_{\text{Residual Matrix }R}.$ (8b)
Lemma 1: The residual matrix $R$ is negligible compared to the naive update $\hat{\theta}_{i}^{\prime}$ in terms of the Frobenius norm.
Proof: Our goal is to show that the error ratio $\left\|R\right\|_{\text{F}}^{2}\,/\,\left\|\hat{\theta}_{i}^{\prime}\right\|_{\text{F}}^{2}$ is small. We will proceed by finding a large enough numerator $\left\|R\right\|_{\text{F}}^{2}$ and small enough $\left\|\hat{\theta}_{i}^{\prime}\right\|_{\text{F}}^{2}$ in the LoRA context and show that the upper bound of the error ratio is small.
Numerator $\left\|R\right\|_{\text{F}}^{2}$ First, we search for conditions that potentially lead to residual matrices with larger Frobenius norm. We apply compact SVD to the patches $\Delta\theta_{i}$ and $\Delta\theta_{i}^{\prime}$, designating subscripts 1 and 2 for SVD matrices, respectively:
$\left\|R\right\|_{\text{F}}^{2}$ $=\left\|\Delta\theta_{i}^{\prime}-\Delta\theta_{i}\right\|_{\text{F}}^{2}$ (9a)
$=\left\|U_{2}\Sigma_{2}V_{2}^{T}-U_{1}\Sigma_{1}V_{1}^{T}\right\|_{\text{F}}^{2}$ (9b)
$=\left\|\sum_{\ell=1}^{r_{2}}\sigma_{\ell}^{(2)}u^{(2)}_{\ell}v^{(2)T}_{\ell}-\sum_{k=1}^{r_{1}}\sigma_{k}^{(1)}u^{(1)}_{k}v^{(1)T}_{k}\right\|_{\text{F}}^{2}.$ (9c)
Here, the typical order of magnitude for $r_{1}$ and $r_{2}$ is about $10$.
1. When there is no intersection between the two pairs of singular vector subspaces, namely, $\mathcal{C}(U_{1})\cap\mathcal{C}(U_{2})=\varnothing$ and $\mathcal{C}(V_{1})\cap\mathcal{C}(V_{2})=\varnothing$, the two terms in (9c) may be combined to form a valid compact SVD of rank $r_{1}+r_{2}$ as follows:
$\left\|R\right\|_{\text{F}}^{2}$ $=\left\|\sum_{\ell=1}^{r_{2}}\sigma_{\ell}^{(2)}u^{(2)}_{\ell}v^{(2)T}_{\ell}+\sum_{k=1}^{r_{1}}\sigma_{k}^{(1)}\cdot(-u^{(1)}_{k})v^{(1)T}_{k}\right\|_{\text{F}}^{2}$ (10a)
$=\left\|\sum_{k^{\prime}=1}^{r_{1}+r_{2}}\sigma_{k^{\prime}}^{(1,2)}u_{k^{\prime}}v_{k^{\prime}}^{T}\right\|_{\text{F}}^{2},$ (10b)
where $\sigma_{1}^{(1,2)},\dots,\sigma_{r_{1}+r_{2}}^{(1,2)}$ is a list of descending ordered positive numbers sampled without replacement from $\{\sigma_{k}^{(1)}\}_{k=1}^{r_{1}}$ and $\{\sigma_{\ell}^{(2)}\}_{\ell=1}^{r_{2}}$. Applying the Frobenius norm property to the SVD representation of a matrix, we obtain
$\left\|R\right\|_{\text{F}}^{2}=\sum_{k^{\prime}=1}^{r_{1}+r_{2}}\left[\sigma_{k^{\prime}}^{(1,2)}\right]^{2}=\sum_{\ell=1}^{r_{2}}\left[\sigma_{\ell}^{(2)}\right]^{2}+\sum_{k=1}^{r_{1}}\left[\sigma_{k}^{(1)}\right]^{2}.$ (11)
This is the case when the two patches $\Delta\theta_{i}$ and $\Delta\theta_{i}^{\prime}$ contain only orthogonal information. This is not very realistic because the two patches were created on the same downstream task $i$ that should lead to some information in common.
2. When the two patches have are oppositely embedded in one of the singular value subspaces, e.g., $U_{2}=-U_{1}$ and $V_{2}=V_{1}$, the two terms in (9c) can be merged and singular values with the same ranking will be summed up, namely,
$\left\|R\right\|_{\text{F}}^{2}$ $=\left\|\sum_{\ell=1}^{r_{2}}\sigma_{\ell}^{(2)}u^{(2)}_{\ell}v^{(2)}_{\ell}+\sum_{k=1}^{r_{1}}\sigma_{k}^{(1)}\cdot(-u^{(1)}_{k})v^{(1)T}_{k}\right\|_{\text{F}}^{2}$ (12a)
$=\left\|\sum_{\ell=1}^{r_{2}}\left[\sigma_{\ell}^{(2)}+\sigma_{\ell}^{(1)}\right]u^{(2)}_{\ell}v^{(2)}_{\ell}\right\|_{\text{F}}^{2}$ (12b)
$=\sum_{\ell=1}^{r_{2}}\left[\sigma_{\ell}^{(2)}+\sigma_{\ell}^{(1)}\right]^{2},$ (12c)
which can be easily shown that it is larger than the orthogonal case (11) due to the extra interaction term $\sum_{\ell=1}^{r_{2}}\sigma_{\ell}^{(2)}\sigma_{\ell}^{(1)}$. When $\sigma_{\ell}^{(2)}=\sigma_{\ell}^{(1)}$, this case corresponds to two patches having exactly opposite gradient update directions, which again, is not very realistic because the same downstream tasks are used to generate the update directions. Equations (11) and (12c) both correspond to extreme conditions, and the continuum in between should be more realistic. We will use the large numerator (12c) to examine the error ratio.
Denominator $\left\|\hat{\theta}_{i}^{\prime}\right\|_{\text{F}}^{2}$ We continue to apply compact SVD to the continued pretrained model $\theta^{\prime}$ and patch $\Delta\theta_{i}$, designating subscripts 0 and 1 for SVD matrices, respectively:
$\left\|\hat{\theta}_{i}^{\prime}\right\|_{\text{F}}^{2}$ $=\left\|\theta^{\prime}+\Delta\theta_{i}\right\|_{\text{F}}^{2}$ (13a)
$=\left\|U_{0}\Sigma_{0}V_{0}^{T}+U_{1}\Sigma_{1}V_{1}^{T}\right\|_{\text{F}}^{2}$ (13b)
$=\left\|\sum_{j=1}^{r_{0}}\sigma_{j}^{(0)}u^{(0)}_{j}v^{(0)T}_{j}+\sum_{k=1}^{r_{1}}\sigma_{k}^{(1)}u^{(1)}_{k}v^{(1)T}_{k}\right\|_{\text{F}}^{2}.$ (13c)
Here, the typical order of magnitude for $r_{0}$ is about $100$.
1. When there is no intersection between the two pairs of singular vector subspaces, the two terms may be combined to form a valid compact SVD of rank $r_{0}+r_{1}$. Hence, similar to (11), we have
$\left\|\hat{\theta}_{i}^{\prime}\right\|_{\text{F}}^{2}=\sum_{j=1}^{r_{0}}\left[\sigma_{j}^{(0)}\right]^{2}+\sum_{k=1}^{r_{1}}\left[\sigma_{k}^{(1)}\right]^{2}.$ (14)
2. When the basis vectors of singular matrices (with one matrix having opposite signs) of the patch can be found in the singular matrices of the continued pretrained model, we are able to combine the two terms in (13c) by using the basis vectors of $U_{0}$ and $V_{0}$ as follows:
$\left\|\hat{\theta}_{i}^{\prime}\right\|_{\text{F}}^{2}$ $=\left\|\sum_{j=1}^{r_{0}}\sigma_{j}^{(0)}u^{(0)}_{j}v^{(0)T}_{j}+\sum_{j=1}^{r_{0}}\sigma_{j}^{(1^{\prime})}u^{(0)}_{j}v^{(0)T}_{j}\right\|_{\text{F}}^{2},$ (15a)
$=\sum_{j=1}^{r_{0}}\left[\sigma_{j}^{(0)}-\sigma_{j}^{(1^{\prime})}\right]^{2},$ (15b)
where we define an auxiliary symbol
$\sigma^{(1^{\prime})}_{j}=\begin{cases}\sigma^{(1)}_{k},&\exists k\in[1,r_{1}]\text{ s.t. }u_{j}^{(0)}=u_{k}^{(1)},\\ 0,&\text{other }k.\end{cases}$ (16)
Both (14) and (15b) correspond to the extreme cases. Since (15b) leads to a smaller denominator, we will use it to examine the error ratio.
Error Ratio. Using (12c) and (15b), a pessimistic error ratio can be approximated and then up bounded as follows:
$\frac{\left\|R\right\|_{\text{F}}^{2}}{\left\|\hat{\theta}_{i}^{\prime}\right\|_{\text{F}}^{2}}$ $\approx\frac{\sum_{\ell=1}^{r_{2}}\left[\sigma_{\ell}^{(2)}+\sigma_{\ell}^{(1)}\right]^{2}}{\sum_{j=1}^{r_{0}}\left[\sigma_{j}^{(0)}-\sigma_{j}^{(1^{\prime})}\right]^{2}}$ (17a)
$\leq\frac{r_{2}\cdot\max_{\ell}\left[\sigma_{\ell}^{(2)}+\sigma_{\ell}^{(1)}\right]^{2}}{r_{0}\cdot\min_{j}\left[\sigma_{j}^{(0)}-\sigma_{j}^{(1^{\prime})}\right]^{2}}$ (17b)
$=\frac{r_{2}}{r_{0}}\cdot\nu,$ (17c)
where $\nu=\max_{\ell}\left[\sigma_{\ell}^{(2)}+\sigma_{\ell}^{(1)}\right]^{2}\Big{/}\min_{j}\left[\sigma_{j}^{(0)}-\sigma_{j}^{(1^{\prime})}\right]^{2}$ is a singular-value based constant. Given that the rank $r_{2}$ of the patch is at least one order of magnitude smaller than the rank $r_{0}$ of the pretrained model, we conclude the residual matrix term is negligible compared to the naive update term.